# Spatial and Spectral Characterization of the Gravitational-wave Background with the PTA Optimal Statistic

### v1

@@ Appendix B Uncertainty on the OS cross-correlation estimators
Unlike with the traditional OS, we cannot simplify this equation any further unless we make an additional assumption. One logical assumption, detailed in Section III, is that of the narrowband-normalized PFOS. This form of the PFOS assumes that all GW power comes only from a single frequency. This leads us to replace $\Phi(f_{k})$ with $\tilde{\phi}(f_{k})$, effectively zeroing the assumed spectral power in other frequencies. This lets us express the narrowband-normalized PFOS’ uncertainty as
@@ Appendix C Covariance of pairwise cross-correlation estimates
$\rho_{ab}=\mathcal{N}_{ab}\,\bm{\delta t}^{T}_{a}\,\bm{P}^{-1}_{a}\,\bm{F}_{a}\,\hat{\phi}\,\bm{F}^{T}_{b}\,\bm{P}^{-1}_{b}\,\bm{\delta t}_{a},$ (71)
@@
In this form, this equation is the same as the one found in Equation $69$ of Section IV.B of Johnson et al. [22]. However, implementing pulsar-pair covariance in this form will be computationally expensive due to the $\bm{Q}_{ab}$ matrix being $(N_{\mathrm{TOA},a}\times N_{\mathrm{TOA},b})$. Instead we can adopt the rank-reduced matrices to compress our matrix representations. It also becomes more helpful to identify the unique cases of $\bm{C}_{ab,cd}$. These three cases are: 00 matching pulsars ($\bm{C}_{ab,cd}$), $1$ matching pulsar ($\bm{C}_{ab,ac}$), and $2$ matching pulsars ($\bm{C}_{ab,ab}$).
@@
In this form, we can cycle our matrices inside the trace and identify the rank-reduced matrix product $\bm{Z}_{a}=\bm{F}^{T}_{a}\bm{P}^{T}_{a}\bm{F}_{a}$. This lets us simplify our equation into

### v2

@@ Appendix B Uncertainty on the OS cross-correlation estimators
Unlike with the traditional OS, we cannot simplify this equation any further unless we make an additional assumption. One logical assumption, detailed in Section III, is that of the narrowband-normalized PFOS. This form of the PFOS assumes that each pulsar measures the GW frequencies independently. This leads us to replace $\Phi(f_{k})$ with $\tilde{\phi}(f_{k})$, making these measurements independent of any assumed spectral shape. This lets us express the narrowband-normalized PFOS’ uncertainty as
@@
While this may seem an attractive alternative, subsection V.2 and subsection V.3 detail how this additional simplification is not valid for realistic data.
@@ Appendix C Covariance of pairwise cross-correlation estimates
$\rho_{ab}=\mathcal{N}_{ab}\,\bm{\delta t}^{T}_{a}\,\bm{P}^{-1}_{a}\,\bm{F}_{a}\,\hat{\phi}\,\bm{F}^{T}_{b}\,\bm{P}^{-1}_{b}\,\bm{\delta t}_{b},$ (71)
@@
In this form, this equation is the same as the one found in Equation $69$ of Section IV.B of Johnson et al. [22]. However, implementing pulsar-pair covariance in this form will be computationally expensive due to the $\bm{Q}_{ab}$ matrix being $(N_{\mathrm{TOA},a}\times N_{\mathrm{TOA},b})$. Instead we can adopt the $\bm{X}$ and $\bm{Z}$ rank-reduced matrices to compress our matrix representations. It also becomes more helpful to identify the unique cases of $\bm{C}_{ab,cd}$. These three cases are: 00 matching pulsars ($\bm{C}_{ab,cd}$), $1$ matching pulsar ($\bm{C}_{ab,ac}$), and $2$ matching pulsars ($\bm{C}_{ab,ab}$).
@@
In this form, we can cycle our matrices inside the trace and identify the rank-reduced matrix product $\bm{Z}_{a}=\bm{F}^{T}_{a}\bm{P}^{-1}_{a}\bm{F}_{a}$. This lets us simplify our equation into